OpenStudy (anonymous):

What is the standard form of the equation of the line through (0, -3) with slope 2/5.

OpenStudy (anonymous):

find the point slope form first using \[y-y_1 = m(x-x_1)\]m = slope. then put it in standard form of \[ax+by = c\]

OpenStudy (anonymous):

y - -3 = 2/5(x-0) can you go from here?

OpenStudy (anonymous):

I did that and the test I need help with doesn't have the answer I came up with.

OpenStudy (anonymous):

I keep getting a negative answer instead of the positive the test has

OpenStudy (anonymous):

show us your answer

OpenStudy (anonymous):

Just redid it!got it ! Thank you!

OpenStudy (anonymous):

No problem